Determine the area of a triangle having the following measurements. B = 83°22, a = 6, and c:8
Answer #1

Solution:

Area of (Side-Angle-Side) triangle = (absinB)/2

B=83^{\circ}22'=83^{\circ}+(22/60)'=83^{\circ}+(22/60)^{\circ}=83.367^{\circ}

\textup{Area of triangle =}\frac{ab\times sinC}{2}=\frac{8\times 6\times sin(83.367^{\circ})}{2}

\textup{Area of triangle =}\frac{48\times 0.9933}{2}=23.84\textup{ sq.units}

\boldsymbol{\textup{Area of triangle = }23.84\textup{ sq.units}}

1596049332255_image.png